OpenStudy (anonymous):

a mixture of 10.0 g of Ne and 10.0 g AR have a total pressure of 1.6 atm. what is the partial pressure of Ne? Actually I know the answer but I didn't know how they get 2/3

OpenStudy (photon336):

Convert both gases to moles: 10 grams Neon X (mol/20 grams) = 0.5 mol neon. 10 grams of Argon X (mol/40grams) 0.25 mol argon. Mol fraction total number of moles (0.25+0.50) = (0.75)moles Argon (0.25/.75)*1.6atm = 0.53 atm Neon. (0.5/.75)*1.6 atm = 1.07 atm